OpenStudy (anonymous):

To open these doors, you must speak three functions in standard form. One function, f(x), with two real rational solutions. One function, g(x), with two real irrational solutions. One function, h(x), with two complex solutions.

OpenStudy (anonymous):

f(x) = x^2 -49 has two real solutions (zeros) x^2 - 49 = 0 (x+7)(x-7) = 0 x = -7 and x =7 --------------------------------------------------------------- f(x) = x^2 -7 has two irrational solutions (zeros) x^2 = 7 x^2 = sqrt (7) and -sqrt(7) ----------------------------------------------------------------- f(x) = x^2 + 1 has two imaginary solutions x^2 + 1 = 0 x^2 = -1 x = i and -i (sqrt(-1) and -sqrt(-1))